solve by system of addition 3x + 5y = -11
x - 2y = 11

Answer by checkley77(12844)

3x + 5y = -11
x - 2y = 11 multiply by -3 & add.
-3x+6y=-33
3x+5y=-11
----------------
11y=-44
y=-44/11
y=-4 ans.
3x+5*-4=-11
3x-20=-11
3x=-11+20
3x=9
x=9/3
x=3 ans.
Proof:
3-2*-4=11
3+8=11
11=11